Maya's test scores are 98, 90, and 86. If he must average at least 90 points to get an A in the class, what must he earn on the next test.
Write and solve each inequality.

Answer by Cromlix(4381)

An average of 90 points for four tests.
Total required for four tests = 4 * 90 = 360
Total attained so far = 98+90+86=274
360 - 274 = 86
Maya must score 86 points in the 4th test
to obtain an average of 90 points overall.
Hope this helps.
:-)